OpenStudy (anonymous):

f(x)=x+6/5. find f^-1(6). the answer is 12/5, correct?

OpenStudy (anonymous):

i got 24/5

OpenStudy (anonymous):

thats not one of the possibilities...

OpenStudy (anonymous):

the answer is one of these: 12/5 5/12 36 -24 24

OpenStudy (anonymous):

is f(x) = x + (6/5) or is it (x+6)/5

OpenStudy (anonymous):

the second

OpenStudy (anonymous):

then its 24

OpenStudy (anonymous):

why

OpenStudy (anonymous):

To find the inverse of f(x), swap the x and y and solve for y \[f(x) = \frac{x+6}{5}\] so \[f^{-1} = 5x - 6\] plug in 6 to get 30-6 = 24

OpenStudy (anonymous):

how did u get 5x-6?

OpenStudy (anonymous):

To find the inverse of f(x), swap the x and y and solve for y

OpenStudy (anonymous):

u said that already -_-

OpenStudy (anonymous):

ie \[x = \frac{y+6}{5}\]

OpenStudy (anonymous):

I know, but that was the answer to your question